A 1000 kg racing car accelerates from rest at a constant rate and covers a distance of 300 m in 10 s. What is the car's acceleration? (in m/s2)
1
Expert's answer
2018-09-17T15:26:08-0400
The distance
s=v_i t+(at^2)/2
Since v_i=0, we get
a=2s/t^2 =(2×200 m)/(〖10〗^2 s^2 )=4 m/s^2